OpenStudy (hlambach):

Find the area of an equilateral triangle with radius 5m. Leave your answer in simplest radical form. Rest in comments.

OpenStudy (hlambach):

A. 225/4 \[\sqrt{3}\] B. 45/4 \[\sqrt{3}\] C. 75/2 \[\sqrt{3}\] D. 15/2 \[\sqrt{3}\]

OpenStudy (hlambach):

Now, I keep getting 25 sqrt (3) I'm guessing the radius is from the center to the bottom like this  then, i would use the triangle area formula a= 1/2 x b x h a= 1/2 x 10 x 5 sqrt(3) a= 5 x 5 sqrt(3) a = 25 sqrt(3)

OpenStudy (hlambach):

But that answer is none of the options... So, am I right or are one of the options actually the answer

OpenStudy (hlambach):

whoops i messed up, "Find the area of an equilateral triangle with radius 5 sqrt(3)m. Leave your answer in simplest radical form. Rest in comments."

OpenStudy (anonymous):

The area of an equilateral triangle can be found with the simple formula \[\frac{ s ^{2}\sqrt{3} }{ 4 } \] s= one side length, and since its equilateral, all sides are the same. 10^2 (sqt3)/4 100(sqt3)/4 - reduce 25sqt3 hmmmm, it appears you're right. i suppose A is a typo?

OpenStudy (hlambach):

hum, i don't know... whats really getting me is what the heck is the radius of a triangle? i might be wrong by assuming its the center to the bottom, it actually might be the side... so a= 5sqrt(3)^2 sqrt(3) / 4 a= 15sqrt(3) / 4 a = 15/4 sqrt(3) but thts also none of the options

ganeshie8 (ganeshie8):

triangle has sides and angles. not radius and arcs.

OpenStudy (hlambach):

exactly, so thats why i am so confused...

ganeshie8 (ganeshie8):

yeah must be question is incorrect. leave this question and move on... Also, call your teacher and inform :)

OpenStudy (hlambach):

OK, thanks for trying to help, anyways. :) Thanks, everyone!

ganeshie8 (ganeshie8):

np :)